In Statistics and Probability 50-item test of Grade 11 ABM E the mean of their scores is 37 and a standard deviation of 8.

       a. Find the lowest score of the upper 20% of the distribution (80th Percentile)

       b. What is the highest score of the 35th Percentile (lower 35% of the distribution)?



1
Expert's answer
2021-10-11T05:13:32-0400

a.

"P=0.8"

"z=0.845"

"z=\frac{x-\mu}{\sigma}"

"x=z\sigma +\mu=0.845\cdot 8+37=43.76"

lowest score of the upper 20% of the distribution = "44"


b.

"P=0.35"

"z=-0.385"

"x=-0.385\cdot 8+37=33.92"

the highest score of the 35th Percentile = 34.